find the supplementary angle of 79 degree
Answer by RAY100(1637)

supplement means two angles equal 180 degrees,, a+b=180
.
if one angle is 79,,,79 +b = 180,,,,,b= 101
.
.
check,,,,,,101 +79 =180,,,,,,ok